find the general value of the equation tan²a+cot²a=2

Grade:11

2 Answers

SRC Jayanth
7 years ago
tan²a+cot²a=2
cot²a+2 = 4tan²a+
tan²a+cot²a + 2tana cota = 4
[tana + cota ]2=4
 
tana + (1/ tana) =2
tan²a + 1 = 2 tana
sec2 a = 2 tana
simplify by writing interms of sin &cos
1=2sina cosa
sin2a = 1
a=45o
 

 
SRC Jayanth
7 years ago
in the second step
i have written wrong
tan²a+cot²a+2 = 4
this is correct :- tan²a+cot²a + 2= 4
i has posted wrong